Alternative Financing Plans

Frey Co. is considering the following alternative financing plans:

Plan 1 Plan 2
Issue 10% bonds (at face value) $1,080,000 $540,000
Issue preferred $1 stock, $10 par 900,000
Issue common stock, $5 par 1,080,000 720,000

Income tax is estimated at 40% of income.

Determine the earnings per share on common stock, assuming that income before bond interest and income tax is $756,000.

Enter answers in dollars and cents, rounding to two decimal places.

Plan 1 $ _____ Earnings per share on common stock
Plan 2 $ _____ Earnings per share on common stock
Answer #1
Plan 1 $ 1.80 Earnings per share on common stock
Plan 2 $ 2.30 Earnings per share on common stock
Workings:
Plan 1 Plan 2
Income before bond interest and income tax 756000 756000
Less: Bond interest expense 108000 54000
Income before income tax 648000 702000
Less: Income tax expense 40% 259200 280800
Net income 388800 421200
Less: Preferred Dividends 0 90000
Net income for Common Stockholders 388800 331200
Divide by Common shares outstanding 216000 144000
Earnings per share on common stock 1.80 2.30
Common shares outstanding:
Plan 1 = 1080000/5 = 216000
Plan 2 = 720000/5 = 144000